Complete the balance sheet and sales information using the following financial data:

Total assets turnover: 1.5x
Days sales outstanding: 35 daysa
Inventory turnover ratio: 7x
Fixed assets turnover: 3.5x
Current ratio: 2.1x
Gross profit margin on sales: (Sales - Cost of goods sold)/Sales = 25%
aCalculation is based on a 365-day year. Do not round intermediate calculations. Round your answer to the nearest cent.

Balance Sheet
Cash $   Current liabilities $  
Accounts receivable    Long-term debt 100,000
Inventories    Common stock   
Fixed assets    Retained earnings 120,000
Total assets $400,000 Total liabilities and equity $  
Sales $   Cost of goods sold $  
Answer #1

Total Asset Turnover = 1.5

Sales/Total Assets = 1.5

Sales = 1.5*400,000 = $600,000

Days Sales Outstanding = 35

365*Accounts receivables/Sales = 35

Accounts Receivables = 35*600,000/365

= $57,534.25

COGS = 600,000*75% = $450,000

Inventory Turnover Ratio = 7

COGS/Inventory = 7

Inventory = 450,000/7

= $64,285.71

Fixed assets = 600,000/3.5 = $171,428.57

Balance Sheet

Cash

$106,751

Current liabilities

$108,843.54

Accounts receivable

57534.25

Long-term debt

100,000

Inventories

64285.71

Common stock

$71,156.46

Fixed assets

171428.6

Retained earnings

120,000

Total assets

$400,000

Total liabilities and equity

$400,000

Sales

600000

Cost of goods sold

450000